# Fan distributions via Tverberg partitions and Gale duality

### v1

@@ 6.2. Test Space
Our test space will be the direct sum of two $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-modules. First, consider the regular representation $\mathbb{R}[\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}]=\{\sum_{(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}}r_{(i,j)}\cdot(i,j)\mid r_{(i,j)}\in\mathbb{R}\,\,\text{for all}\,\,(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}\}$, for which the $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-action is given by setting $(k,\ell)\cdot(i,j)=(k+i,\ell+j)$ for each $(k,\ell),(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$ and extending linearly. Inside the regular reperesentation is the subrepresentation $U_{r}=\{\sum_{(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}}r_{(i,j)}\cdot(i,j)\mid\sum_{(i,j)}r_{i,j}=0\}$ consisting of all elements of the regular representation whose coefficients sum to zero. In other words, $U_{r}$ is the complement of the diagonal 1-dimensional trivial subrepresentation in $\mathbb{R}[\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}]$ and so has dimension $r^{2}-1$. For each $m\geq 1$, we then let $U_{r}^{\oplus m}$ be the $m$-fold direct sum of $U_{r}$, which we equip with the diagonal $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-action.

### v2

@@ 6.2. Test Space
Our test space will be the direct sum of two $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-modules. First, consider the regular representation $\mathbb{R}[\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}]=\{\sum_{(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}}r_{(i,j)}\cdot(i,j)\mid r_{(i,j)}\in\mathbb{R}\,\,\text{for all}\,\,(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}\}$, for which the $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-action is given by setting $(k,\ell)\cdot(i,j)=(k+i,\ell+j)$ for each $(k,\ell),(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$ and extending linearly. Inside the regular representation is the subrepresentation $U_{r}=\{\sum_{(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}}r_{(i,j)}\cdot(i,j)\mid\sum_{(i,j)}r_{i,j}=0\}$ consisting of all elements of the regular representation whose coefficients sum to zero. In other words, $U_{r}$ is the complement of the diagonal 1-dimensional trivial subrepresentation in $\mathbb{R}[\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}]$ and so has dimension $r^{2}-1$. For each $m\geq 1$, we then let $U_{r}^{\oplus m}$ be the $m$-fold direct sum of $U_{r}$, which we equip with the diagonal $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-action.